OpenStudy (anonymous):

What is the slope of the line represented by (1/8)x + 3y=3

OpenStudy (anonymous):

1/8x + 3y=3 3y = 1/8x + 3 You divide straight through by 3 to get your answer

OpenStudy (anonymous):

it is easier when you have the equation in standard form

OpenStudy (anonymous):

y = mx+b

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

your slope is going to be m and b is the intercept

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

So would the answer be 2.6

OpenStudy (anonymous):

no... the slope is represented by m in the equation so m = 1/8

OpenStudy (anonymous):

do you see that?

OpenStudy (anonymous):

So once you put it in standard form you get 3y=1/8x+3 right? Then you divide by 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now y=.04167+1? What am I missing?